# Picocavity-enhanced Raman Spectroscopy of Physisorbed H2 and D2 Molecules

### v1

@@ Acknowledgements.
To verify the contribution of LSPR to the TERS intensity, we examined the excitation wavelength () dependence. Figure S3(a) shows a STML spectrum recorded over a clean Ag(111) surface with a Ag tip. The LSPR results in a broad spectral feature in the visible range with a maximum at 560 nm. After recording the STML, we exposed the surface to the D2 gas, conducting TERS measurements with = 532 and 633 nm [Figs. S3(b) and S3(c), respectively]. The peak intensity of each mode significantly changes with , whereas the Raman shift value is unchanged. The vibrational intensity is weakened with = 633 nm, indicating the corresponding scattering light is not efficiently enhanced by the LSPR because the peak position ( = 779 nm) is located in the weak resonance range of the LSPR [Fig. S3(a)]. In contrast, both the rotational and vibrational peaks become intense with = 532 nm because they are more resonant with the LSPR.

### v2

@@ Acknowledgements.
To verify the contribution of LSPR to the TERS intensity, we examined the excitation wavelength () dependence. Figure S3(a) shows a STML spectrum recorded over a clean Ag(111) surface with an Ag tip. The LSPR results in a broad spectral feature in the visible range with a maximum at 560 nm. After recording the STML, we exposed the surface to the D2 gas, conducting TERS measurements with = 532 and 633 nm [Figs. S3(b) and S3(c), respectively]. The peak intensity of each mode significantly changes with , whereas the Raman shift value is unchanged. The vibrational intensity is weakened with = 633 nm, indicating the corresponding Raman scattering is not efficiently enhanced by the LSPR because the peak position ( = 779 nm) is located in the weak resonance range of the LSPR [Fig. S3(a)]. In contrast, both the rotational and vibrational peaks become intense with = 532 nm because they are more resonant with the LSPR.
II.3 C. Tip-Height Dependence of Peak Intensity and Width
To obtain the waterfall plots shown in Fig. 3 of the main text, we conducted tip approach-and-retraction procedures while recording the TERS spectra [54, 41]. Figure S10(a) is the full dataset for the H2 “vib” mode shown in Fig. 3(b) of the main text. During recording TERS spectra, the tip height was lowered stepwise and then the tip was retracted from the surface. The symmetrical waterfall plot with respect to the closest tip height indicates the non-destructive measurement in the tip approach-and-retraction process. Figure S10(b) shows the horizontal line profiles of the waterfall plot for the H2 “vib” mode at the points indicated by the arrows in the same color. The three spectra indicates the red-shift and broadening of the H2 “vib” peak as the gap distance decreases.
During recording the waterfall TERS plot, we occasionally observed a sudden intensity change caused presumably by the modification of the tip apex structure due to the strong tip–surface interaction at the small gap. Figure S10(c) shows an example, where a sudden drop of the TERS intensity was observed. The modification of the plasmon background intensity implies that the atomic-scale structure was changed by the strong attractive force exerted when the tip was brought close to the surface. At the same time as the plasmon feature, the D2 “rot” peak was also dropped, while the frequency of the “rot” mode is unchanged. This observation is in good agreement with the enhancement mechanism described in the main text (Fig. 2); the picocavity-derived electric field plays a crucial role in the TERS intensity of physisorbed hydrogen molecules.
To determine the experimental peak position shift as a function of the tip height in Fig. 4(e) of the main text, the line profile of the waterfall plots at each tip height [Fig. 3 in the main text; see also Fig. S10(b)] was fitted by a Gaussian peak with a constant background. The top and bottom panels of Fig. S11 shows the peak position and width, respectively, as a function of the tip height for the rotational (a) and vibrational (b) modes. The error bars indicate the standard deviation of the fitting. As the tip approaches, the peak intensity increases, reducing the error. The point at which the error becomes below 2 cm-1 by the tip approach was used as the reference point [the arrows in the top panel of Fig. S11(a) and (b)] and the peak position at this reference point was set to the origin of the peak shift (0 cm-1 of the top panel). For the H2 and D2 “vib” modes [top panel of Fig. S11(b)], the plots after the reference points were used as the experimental plots shown in Fig. 4(e) of the main text. In Fig. 4(e) of the main text, the experimental tip height pm is aligned to the tip–sample distance nm for the calculations. Although the accurate determination of the experimental tip–sample gap distance is difficult, this alignment is reasonable; as described in Sect. I-E, pm is comparable with pm and nm.
As shown in Fig. S11(a), the frequency and width of the H2 and D2 “rot” peaks are almost unchanged by the tip. In contrast, the bottom panel Fig. S11(b) shows the strong peak broadening for the H2 “vib” mode [see also Fig. S10(b)] whereas the peak width for D2 “vib” is almost unchanged. A possible origin of the peak broadening is the anharmonicity, which can modify the lifetime of molecular vibrations [71]. It is also entirely possible that if neighboring molecules influence the experimental signal [see Fig. 2(c) in the main text], they would lead to the high-frequency shoulder of the vibration transition for H2, as these molecules would still lie on the surface [Fig. 3(c)] and would therefore present a higher than the ones on the tip, which instead provide the most enhanced signal; however, this effect may not be pronounced because the coverage of is less than that of D2 (see Sect. I-C). Another possible origin if the movement of the molecule in the junction; because the double-well potential barrier is very small [see Fig. 4(d) in the main text], the H2 molecule near the tip apex can move to a metastable site near the surface or adjacent adsorption sites on the surface [18] by a small voltage bias or plasmonic field applied in the STM junction. This movement instantaneously occurring during the acquisition time of the Raman spectrum can also modify the peak shape of the vibrational mode. We expect that the coexistence of these factors leads to the characteristic peak width depending on the tip height.